OpenStudy (anonymous):

Simplify 4x + 5x - 5y + 3y 9x + 2y 10x + 2y , 9x - 2y, 10x + 3y , or 9x - 3y

OpenStudy (aroub):

Combine like terms!

OpenStudy (aroub):

=> (4x+5x) + (-5y+3y) = 9x-2y

OpenStudy (anonymous):

is it 9x + 2y

OpenStudy (eyust707):

close!

OpenStudy (eyust707):

whats -5 + 3 = ??

OpenStudy (anonymous):

-2

OpenStudy (eyust707):

do you see where the -2 should go?

OpenStudy (anonymous):

its the answer -5+3=-2